OpenStudy (anonymous):

find dp/dq p=5+ (1/cotq) im getting csc^2q/cot^2q but the book says sec^2q, and theyre right i need help!! wa did i do wrong?

OpenStudy (anonymous):

Simplfy the trig expressions, I think they are equivalent.\[\csc ^{2}q/\cot ^{2}q=1/\sin ^{2}q*\sin ^{2} q/\cos ^{2}q=1/\cos ^{2}q=\sec ^{2}q\]

OpenStudy (amistre64):

my eyes!!! my precious little eyes!!!

OpenStudy (anonymous):

so sec^2q can equal 1/cos^2q? theres no rule that says i cant manipulate these figures like this?

OpenStudy (anonymous):

cuz i knew that secq=1/cosq i just didnt knw i could manipulate em like that

OpenStudy (anonymous):

That's the definition of sec (reciprocal of cos).